OpenStudy (anonymous):

Evaluate the summation of 3 n plus 2, from n equals 1 to 14..

OpenStudy (ankit042):

3n+2 from n 1 to 14 3(sum(n))+(14)*2

OpenStudy (ankit042):

sum(x) from 1 to n = n(n+1)/2

OpenStudy (ankit042):

so it becomes 3(14(14+1)/2) +28 @KaylaHake can you solve now?

OpenStudy (anonymous):

yeah i was able to solve it

OpenStudy (anonymous):

could you help me with another one?

OpenStudy (anonymous):

Evaluate the summation of 25 times 0.3 to the n plus 1 power, from n equals 2 to 10..

OpenStudy (ankit042):

sure post a new question and tag me in

OpenStudy (anonymous):

@ankit042

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 3n+2 }\) is, as you can tell, an arithmetic sequence where the rate of change or the growth of each next term is 3. (you can call it a "slope" 3) \(\LARGE\color{blue}{ \bf 3(1)+2=3+2=5 }\) \(\LARGE\color{blue}{ \bf 3(2)+2=6+2=8 }\) \(\LARGE\color{blue}{ \bf 3(3)+2=9+2=11 }\) The formula of a sum for this sequence would be, \(\LARGE\color{blue}{ \bf \frac{14(a_1+a_{14})}{2} }\)

OpenStudy (solomonzelman):

you can find a14 by simply plugging in 14 for n, into 3n+2.

OpenStudy (anonymous):

OPEN THE PIC :)